# GDBA Revisited: Unleashing the Power of Guided Local Search for Distributed Constraint Optimization

### v1

@@ Abstract
Local search is an important class of incomplete algorithms for solving Distributed Constraint Optimization Problems (DCOPs) but it often converges to poor local optima. While GDBA provides a comprehensive rule set to escape premature convergence, its empirical benefits remain marginal on general-valued problems. In this work, we systematically examine GDBA and identify three factors that potentially lead to its inferior performance, i.e., over-aggressive constraint violation conditions, unbounded penalty accumulation, and uncoordinated penalty updates. To address these issues, we propose Distributed Guided Local Search (DGLS), a novel GLS framework for DCOPs that incorporates an adaptive violation condition to selectively penalize constraints with high cost, a penalty evaporation mechanism to control the magnitude of penalization, and a synchronization scheme for coordinated penalty updates. We theoretically show that the penalty values are bounded, and agents play a potential game in our DGLS. Our extensive empirical results on various standard benchmarks demonstrate the great superiority of DGLS over state-of-the-art baselines. Particularly, compared to Damped Max-sum with high damping factors (e.g., 0.7 or 0.9), our DGLS achieves competitive performance on general-valued problems, and outperforms it by significant margins (3.77%–66.3%) on structured problems in terms of anytime results.
@@ 1 Introduction
Distributed Constraint Optimization Problems (DCOPs) (Modi et al., 2005; Fioretto et al., 2018) are a fundamental formalism for cooperative multi-agent systems where a set of autonomous agents coordinate with each other to pursue a global objective via localized communication. DCOPs have been successfully applied to model various real-world applications, including scheduling (Pertzovsky et al., 2024), resource allocation (Monteiro et al., 2012), and smart grids (Fioretto et al., 2017).
@@
Efficiently solving DCOPs remains a long-standing challenge due to the inherent NP-hardness. Over the past decades, many DCOP algorithms have been proposed and are generally categorized into complete algorithms and incomplete algorithms, according to whether they guarantee finding the optimal solutions. Complete algorithms include distributed backtracking search (Modi et al., 2005; Chechetka and Sycara, 2006; Yeoh et al., 2010; Netzer et al., 2012; Litov and Meisels, 2017; Deng et al., 2019) and inference (Petcu and Faltings, 2005, 2007; Chen et al., 2020), which exhaust the solution space by either branch-and-bound or bucket elimination (Dechter, 1998). However, the coordination overheads of these algorithms scale exponentially w.r.t. the problem size (e.g., number of agents, induced width), making complete algorithms unsuitable for large-scale applications.
@@
On the other hand, incomplete algorithms trade the optimality for practical computational overheads (Farinelli et al., 2008; Cohen et al., 2020; Chen et al., 2018; Nguyen et al., 2019; Ottens et al., 2017). Among them, local search (Zhang et al., 2005; Maheswaran et al., 2004a; Pearce and Tambe, 2007; Zivan et al., 2014; Hoang et al., 2018) is an important class of incomplete algorithms, which iteratively refines the solution via local moves. However, these algorithms often prematurely converge to poor local optima due to their greedy nature. As an instantiation of Guided Local Search (GLS) (Voudouris et al., 2010), GDBA (Okamoto et al., 2016) was introduced to provide a comprehensive rule set for breaking out of the local optima by adapting the notion of constraint violation and the manner of cost increase in DBA (Hirayama and Yokoo, 2005). However, the empirical benefits of the strongest GDBA variant ($M,NM,T$) often appear to be marginal compared to well-established baselines like DSA (Zhang et al., 2005) on general-valued problems.
@@
To understand why GDBA underperforms on general-valued problems, we first conduct a pilot study (cf. Figure 1) analyzing the penalty dynamics of GDBA on both random DCOPs and structured problems (e.g., meeting scheduling). Our observations reveal that, on general-valued instances, GDBA uniformly accumulates penalties for nearly all constraints, making each constraint receive heavy but similar attention, which essentially offsets the benefit of breakout. Such ineffective penalization stems from over-aggressive constraint violation conditions (e.g., $NM$) that classify most constraints as violated, monotonic penalty increases that lead to unbounded penalty accumulation, and uncoordinated penalty updates that cause agents to optimize misaligned objective functions. In light of this, we present a novel Distributed Guided Local Search (DGLS) framework for DCOPs, which incorporates an adaptive constraint violation condition based on the costs of each constraint, an evaporation mechanism to avoid unbounded penalty accumulation, and a synchronization scheme to enforce coordinated penalty updates. Specifically, our contributions are:
@@
  • We present a novel DGLS framework that enables efficient GLS for DCOPs. Besides the combinations of different manners and scopes like GDBA, our DGLS also incorporates an adaptive constraint violation condition, an evaporation mechanism, and a penalty synchronization scheme to fully unleash the performance of GLS in solving DCOPs. We also theoretically show that the penalty values are bounded, and agents play a potential game in our DGLS.
  • We compare our DGLS with state-of-the-art DCOP algorithms on various standard benchmarks. Our extensive empirical results show the great potential of DGLS: on general-valued uniform problems and scale-free network problems, DGLS is able to match or perform slightly better than Damped Max-sum (DMS) (Cohen et al., 2020) with a high damping factor (i.e., 0.7 or 0.9), while DGLS outperforms DMS by significant margins on structured problems including 2D lattices (by 3.77%–6.03%), meeting scheduling (by 5.47%–9.45%) and weighted graph coloring (by 61.24%–66.30%) in terms of anytime performance (Zilberstein, 1996; Zivan et al., 2014).
@@ 2.1 Distributed Constraint Optimization Problems
A DCOP (Modi et al., 2005) can be formalized as a tuple $\langle I,X,D,F\rangle$ where $I=\{1,\dots,|I|\}$ is the set of agents, $X=\{x_{1},\dots,x_{|X|}\}$ is the set of variables, $D=\{D_{1},\dots,D_{|X|}\}$ is the set of discrete domains, and $F=\{f_{1},\dots,f_{|F|}\}$ is the set of constraint functions. Each variable $x_{i}$ takes a value from its domain $D_{i}$, and each constraint function $f_{i}:D_{i_{1}}\times\cdots\times D_{i_{k}}\rightarrow\mathbb{R}_{\geq 0}$ defines a cost for each possible combination of variables $scp(f_{i})=(x_{i_{1}},\dots,x_{i_{k}})$. The objective is to find a solution $\tau^{*}=\left(d_{1}^{*},\dots,d_{|X|}^{*}\right)$ such that the total cost is minimized:
@@ 2.2 Guided Local Search
GLS (Voudouris et al., 2010) is a metaheuristic for helping local search algorithms to escape local optima using a penalty system. Specifically, GLS considers the following augmented objective function:
@@
where $f(\cdot)$ is the original objective function, $p_{i}$ is the penalty associated with feature $i$, $\mathbb{I}$ is the indicator function, and $\lambda>0$ is the weight to balance the original objective and penalty. Here, the features are problem-specific. For example, a feature could be an edge from city A to city B in a Traveling Salesman Problem (TSP), or whether a hard clause is satisfied in Boolean Satisfiability (SAT) (Cai and Lei, 2020). When local search converges to a local optimum, GLS will select a subset of features presented in the incumbent solution to increase the associated penalty value, so as to force local search to explore novel solutions. In other words, GLS breaks out of the local optimum by modifying the problem’s objective landscape.
@@ 2.3 Generalized Distributed Breakout Algorithm
GDBA (Okamoto et al., 2016) provides a comprehensive set of rules for breaking out of local optima in DCOPs, which adapts the concepts of constraint violation and cost increase in DBA (Hirayama and Yokoo, 2005) for solving Distributed Constraint Satisfaction Problems (DisCSPs) (Yokoo et al., 1998). Essentially, GDBA can be viewed as an instantiation of GLS where the features can be a full constraint function, a specific row or column within it, or a single cost cell, depending on the scope of changes to the penalty values during breakouts. Specifically, when a Quasi Local Minimum (QLM) (Hirayama and Yokoo, 2005) is detected, each agent in QLM first identifies a set of violated constraints based on the cost values under the incumbent local solution, then independently increases the penalty associated with the features in these violated constraints by 1, which corresponds to selective penalization in GLS. Besides additive penalty like Eq. (2), GDBA’s variants also consider the multiplicative penalty:
@@ 3.1 Motivation
To analyze the dynamics of penalty values of GDBA, we consider its widely used variant $\langle M,NM,T\rangle$ on both sparse random DCOPs and meeting scheduling (Zivan et al., 2014; Maheswaran et al., 2004b) where GDBA demonstrates notably inferior and superior performance, respectively (cf. Section 4). Figure 1(a) plots the mean penalty values in all cost modifiers against iterations.
@@
Finally, agents in QLM independently increase the cost modifiers, which may cause agents to optimize different objectives. Consider a constraint $f_{ij}$ being flagged as violated for the first time. If agent $i$ is in QLM while agent $j$ doesn’t, then the cost modifier for $f_{ij}$ becomes 1 from agent $i$’s side but remains 0 from agent $j$’s side after breakout. Such mismatch introduces asymmetry in cost structures and potentially breaks the correspondence between pure strategy Nash Equilibrium and local optimum (Chapman et al., 2011; Grinshpoun et al., 2013).
@@ 3.2 DGLS Framework
Instead of using fixed rules in GDBA, our DGLS incorporates an adaptive rule for identifying violated constraints (cf. Algorithm 3). Intuitively, for each constraint $f_{ij}$ we first compute its normalized cost $\eta$, where $\check{f}_{ij}=\min_{d_{i},d_{j}}f_{ij}(d_{i},d_{j})$ and $\hat{f}_{ij}=\max_{d_{i},d_{j}}f_{ij}(d_{i},d_{j})$. Then we stochastically mark the constraint as violated with a probability of $\eta$. This also generalizes the utility-based penalization in classical GLS (Voudouris et al., 2010), where the features with the highest score are deterministically penalized.
Our adaptive rule has several unique advantages. First, it eliminates the need to tune the constraint violation condition in GDBA (e.g., NZ,NMNZ,\;NM and $MX$). Second, it perfectly aligns with objective (1) by measuring the “badness" of a constraint with the normalized cost. Particularly, if the constraint cost equals the minimum value, then the constraint cannot be flagged as violated; conversely, if the constraint cost attains the maximum value, then there must be a constraint violation. By selectively penalizing constraints, we avoid over-penalizing the constraints with the cost close to their minimum value, which directs local search to pay more attention to the constraints that incur high costs.
@@ 3.3 Theoretical Results
At the beginning of each round, the cost modifier of the constraint between agent $i$ and agent $j$ from $i$’s side is the same as the counterpart from $j$’s side. [^1] Note that in our algorithm, the first dimension of a table is always for the current agent and the second one is for the neighbor agent, e.g., $f_{ij}=f_{ji}^{T}$. Therefore, by saying “the same as”, we essentially refer to $M_{ij}=M_{ji}^{T}$.
Proof.
We prove the lemma by induction. The lemma holds trivially for the first round where all cost modifiers are initialized to 0. Assume that the lemma holds for round $k$. Then the cost modifiers are same after evaporation since we are using the same discounted factor $\gamma$ for both agent $i$ and $j$. We now consider the following cases for coordinated penalty update (cf. Algorithm 4):
  • Case 1: neither $i$ penalizes $f_{ij}$ nor $j$ penalizes $f_{ji}$. In this case, no entry of the cost modifiers changes and the lemma naturally holds for the $k+1$ round.
  • Case 2: only $i$ penalizes $f_{ij}$. Before performing coordinated penalty update, $j\in\bar{P}_{i}$ and $i\in\tilde{P}_{j}$ (cf. line 17-19 of Algorithm 1). Note that $j\notin\tilde{P}_{i}$ and $i\notin\bar{P}_{j}$. We consider each update scope:
    • Cell scope: agent $i$ updates $M^{(k+1)}_{ij}(d_{i},d_{j})\leftarrow M^{(k)}_{ij}(d_{i},d_{j})+1$ since $j\in\bar{P}_{i}$. Agent $j$ updates $M^{(k+1)}_{ji}(d_{j},d_{i})\leftarrow M^{(k)}_{ji}(d_{j},d_{i})+1$ since $i\in\tilde{P}_{j}$. Given the induction assumption of $M^{(k)}_{ij}=\left(M^{(k)}_{ji}\right)^{T}$, $M^{(k+1)}_{ij}(d_{i},d_{j})=M^{(k+1)}_{ji}(d_{j},d_{i})$ and therefore the lemma holds after update.
    • Table scope: agent $i$ increments all entries $M^{(k+1)}_{ij}(d_{i}^{\prime},d_{j}^{\prime})\leftarrow M^{(k)}_{ij}(d_{i}^{\prime},d_{j}^{\prime})+1$, while agent $j$ correspondingly updates $M^{(k+1)}_{ji}(d_{j}^{\prime},d_{i}^{\prime})\leftarrow M^{(k)}_{ji}(d_{j}^{\prime},d_{i}^{\prime})+1$ for all $d_{i}^{\prime}\in D_{i},d_{j}^{\prime}\in D_{j}$. Given $M^{(k)}_{ij}=\left(M^{(k)}_{ji}\right)^{T}$, it is the case that $M^{(k+1)}_{ij}(d_{i}^{\prime},d_{j}^{\prime})=M^{(k+1)}_{ji}(d_{j}^{\prime},d_{i}^{\prime}),\;\forall d_{i}^{\prime}\in D_{i},d_{j}^{\prime}\in D_{j}$, which preserves symmetry for round $k+1$.
    • Row scope: agent $i$ updates $M^{(k+1)}_{ij}(d_{i},d_{j}^{\prime})\leftarrow M^{(k)}_{ij}(d_{i},d_{j}^{\prime})+1$ while agent $j$ updates $M^{(k+1)}_{ji}(d_{j}^{\prime},d_{i})\leftarrow M^{(k)}_{ji}(d_{j}^{\prime},d_{i})+1$, which preserves symmetry for round $k+1$ given the assumption $M^{(k)}_{ij}(d_{i},d_{j}^{\prime})=M^{(k)}_{ji}(d_{j}^{\prime},d_{i}),\;\forall d_{j}^{\prime}\in D_{j}$.
    • Column scope: the analysis is similar to the row scope case with swapped row and column operation.
  • Case 3: only $j$ penalizes $f_{ji}$. By symmetry with Case 2, where $i\in\bar{P}_{j}$ and $j\in\tilde{P}_{i}$, but $i\notin\tilde{P}_{j}$ and $j\notin\bar{P}_{i}$. The analysis follows the same pattern as Case 2, with roles of $i$ and $j$ exchanged.
  • Case 4: both $i$ penalizes $f_{ij}$ and $j$ penalizes $f_{ji}$. In this case, $j\in\bar{P}_{i}$, $i\in\tilde{P}_{j}$, $j\in\tilde{P}_{i}$, and $i\in\bar{P}_{j}$. We consider each update scope:
    • Cell and table scope: both agents increment the cost modifiers as in Case 2, which preserves the symmetry.
    • Row scope: for each $d^{\prime}_{i}\in D_{i}$, both agent increment the related entries as:
      $M^{(k+1)}_{ij}(d_{i}^{\prime},d_{j})\leftarrow M^{(k)}_{ij}(d_{i}^{\prime},d_{j})+1\quad(\text{agent }i),$
      $M^{(k+1)}_{ji}(d_{j},d_{i}^{\prime})\leftarrow M^{(k)}_{ji}(d_{j},d_{i}^{\prime})+1\quad(\text{agent }j).$
      Similarly, for each $d^{\prime}_{j}\in D_{j}$, we have:
      $M^{(k+1)}_{ij}(d_{i},d_{j}^{\prime})\leftarrow M^{(k)}_{ij}(d_{i},d_{j}^{\prime})+1\quad(\text{agent }i),$
      $M^{(k+1)}_{ji}(d_{j}^{\prime},d_{i})\leftarrow M^{(k)}_{ji}(d_{j}^{\prime},d_{i})+1\quad(\text{agent }j).$
      This maintains symmetry given $M^{(k)}_{ij}=\left(M^{(k)}_{ji}\right)^{T}$. Note that for current assignment pair $(d_{i},d_{j})$, both agents subtract the corresponding entries by 1 to avoid double-counting (cf. line 13-14 of Algorithm 4), which does not affect the symmetry.
    • Column scope: the analysis is similar to the row scope case with swapped row and column operation.
In all cases, the symmetry $M^{(k+1)}_{ij}=\left(M^{(k+1)}_{ji}\right)^{T}$ is preserved, completing the induction step. Therefore, the lemma holds for all rounds ∎
@@
$\textsc{EffCost}_{A}^{(k)}(d_{i},j,d_{j})=1+M_{ij}(d_{i},d_{j})$
$\textsc{EffCost}_{M}^{(k)}(d_{i},j,d_{j})=1\cdot\left[1+M_{ij}(d_{i},d_{j})\right]=1+M_{ij}(d_{i},d_{j}).$
(a) Sparse problems
(b) Dense problems
Figure 2: Performance on random DCOPs
@@
$(A,\gamma,tab)$ is equivalent to MGM for any $\gamma$.
@@
This exactly matches the decision rule for agents in MGM (Maheswaran et al., 2004a). Besides,
$\Delta_{i}$ $=\sum_{j\in\mathcal{N}_{i}}\textnormal{{EffCost}}(d_{i},j,d_{j})-\textnormal{{EffCost}}(d_{i}^{*},j,d_{j})$
$=\sum_{j\in\mathcal{N}_{i}}f_{ij}(d_{i},d_{j})+M_{ij}(d_{i},d_{j})-f_{ij}(d_{i}^{*},d_{j})-M_{ij}(d_{i}^{*},d_{j})$
$=\sum_{j\in\mathcal{N}_{i}}f_{ij}(d_{i},d_{j})-f_{ij}(d_{i}^{*},d_{j})+c_{ij}-c_{ij}$
$=\sum_{j\in\mathcal{N}_{i}}f_{ij}(d_{i},d_{j})-f_{ij}(d_{i}^{*},d_{j}),$

### v2

@@ Abstract
Local search is an important class of incomplete algorithms for solving Distributed Constraint Optimization Problems (DCOPs) but it often converges to poor local optima. While Generalized Distributed Breakout Algorithm (GDBA) provides a comprehensive rule set to escape premature convergence, its empirical benefits remain marginal on general-valued problems. In this work, we systematically examine GDBA and identify three factors that potentially lead to its inferior performance, i.e., over-aggressive constraint violation conditions, unbounded penalty accumulation, and uncoordinated penalty updates. To address these issues, we propose Distributed Guided Local Search (DGLS), a novel GLS framework for DCOPs that incorporates an adaptive violation condition to selectively penalize constraints with high cost, a penalty evaporation mechanism to control the magnitude of penalization, and a synchronization scheme for coordinated penalty updates. We theoretically show that the penalty values are bounded, and agents play a potential game in DGLS. Extensive empirical results on various benchmarks demonstrate the great superiority of DGLS over state-of-the-art baselines. Compared to Damped Max-sum with high damping factors, our DGLS achieves competitive performance on general-valued problems, and outperforms by significant margins on structured problems in terms of anytime results.
Codehttps://github.com/ycdeng-ntu/DGLS
Extended versionhttps://arxiv.org/pdf/2508.06899
@@ 1 Introduction
Distributed Constraint Optimization Problems (DCOPs) (Modi et al. 2005; Fioretto et al. 2018) are a fundamental formalism for cooperative multi-agent systems where a set of autonomous agents coordinate with each other to pursue a global objective via localized communication. DCOPs have been successfully applied to model various real-world applications, including scheduling (Pertzovsky et al. 2024), resource allocation (Monteiro et al. 2012), and smart grids (Fioretto et al. 2017).
@@
Efficiently solving DCOPs remains a long-standing challenge due to the inherent NP-hardness. Over the past decades, many DCOP algorithms have been proposed and are generally categorized into complete algorithms and incomplete algorithms, according to whether they guarantee finding the optimal solutions. Complete algorithms include distributed backtracking search (Modi et al. 2005; Chechetka and Sycara 2006; Yeoh et al. 2010; Netzer et al. 2012; Litov and Meisels 2017; Deng et al. 2019) and inference (Petcu and Faltings 2005, 2007; Chen et al. 2020), which exhaust the solution space by either branch-and-bound or bucket elimination (Dechter 1998). However, the coordination overheads of these algorithms scale exponentially w.r.t. the problem size, making them unsuitable for large-scale applications.
@@
On the other hand, incomplete algorithms trade the optimality for practical computational overheads (Farinelli et al. 2008; Cohen et al. 2020; Chen et al. 2018; Nguyen et al. 2019; Ottens et al. 2017). Among them, local search (Zhang et al. 2005; Maheswaran et al. 2004a; Pearce and Tambe 2007; Zivan et al. 2014; Hoang et al. 2018) is an important class of incomplete algorithms, which iteratively refines the solution via local moves. However, these algorithms often prematurely converge to poor local optima due to their greedy nature. As an instantiation of Guided Local Search (GLS) (Voudouris et al. 2010), Generalized Distributed Breakout Algorithm (GDBA) (Okamoto et al. 2016) was introduced to provide a comprehensive rule set for breaking out of the local optima by adapting the notion of constraint violation and cost increase in DBA (Hirayama and Yokoo 2005). However, the empirical benefits of GDBA often appear to be marginal compared to well-established baselines like DSA (Zhang et al. 2005) on general-valued problems.
@@
To understand why GDBA underperforms on general-valued problems, we first conduct a pilot study (cf. Figure 1) analyzing the penalty dynamics of GDBA on both random DCOPs and structured problems. Our observations reveal that, on general-valued instances, GDBA uniformly accumulates penalties for nearly all constraints, making them receive heavy but similar attention, which offsets the benefit of breakout. Such ineffective penalization stems from over-aggressive constraint violation conditions (e.g., $NM$) that classify most constraints as violated, monotonic penalty increases that lead to unbounded penalty accumulation, and uncoordinated penalty updates that cause agents to optimize misaligned objective functions. In light of this, we present a novel Distributed Guided Local Search (DGLS) framework for DCOPs, which incorporates an adaptive constraint violation condition based on the costs of each constraint, an evaporation mechanism to avoid unbounded penalty accumulation, and a synchronization scheme to enforce coordinated penalty updates. Specifically, our contributions are:
@@
  • We present a novel DGLS framework that enables efficient GLS for DCOPs. Our DGLS incorporates an adaptive constraint violation condition, an evaporation mechanism, and a penalty synchronization scheme to fully unleash the performance of GLS in solving DCOPs. We also theoretically show that the penalty values are bounded, and agents play a potential game in our DGLS.
  • We compare our DGLS with state-of-the-art DCOP algorithms on various standard benchmarks. Our extensive empirical results show the great potential of DGLS: on general-valued problems and scale-free network, DGLS is able to match or perform slightly better than Damped Max-sum (DMS) (Cohen et al. 2020) with a high damping factor, while DGLS outperforms DMS by significant margins on structured problems including 2D lattices, meeting scheduling and weighted graph coloring in terms of anytime performance (Zilberstein 1996; Zivan et al. 2014).
@@ 2.1 Distributed Constraint Optimization Problems
A DCOP (Modi et al. 2005) can be formalized as a tuple $\langle I,X,D,F\rangle$ where $I=\{1,\dots,|I|\}$ is the set of agents, $X=\{x_{1},\dots,x_{|X|}\}$ is the set of variables, $D=\{D_{1},\dots,D_{|X|}\}$ is the set of discrete domains, and $F=\{f_{1},\dots,f_{|F|}\}$ is the set of constraint functions. Each variable $x_{i}$ takes a value from its domain $D_{i}$, and each constraint function $f_{i}:D_{i_{1}}\times\cdots\times D_{i_{k}}\rightarrow\mathbb{R}_{\geq 0}$ defines a cost for each possible combination of variables $scp(f_{i})=(x_{i_{1}},\dots,x_{i_{k}})$. The objective is to find a solution $\tau^{*}=\left(d_{1}^{*},\dots,d_{|X|}^{*}\right)$ such that the total cost is minimized:
@@ 2.2 Guided Local Search
GLS (Voudouris et al. 2010) is a metaheuristic for helping local search algorithms to escape local optima using a penalty system. Specifically, GLS considers the following augmented objective function:
@@
where $f(\cdot)$ is the original objective function, $p_{i}$ is the penalty associated with feature $i$, $\mathbb{I}$ is the indicator function, and $\lambda>0$ is the weight to balance the original objective and penalty. Here, the features are problem-specific. For example, a feature could be an edge from city A to city B in a Traveling Salesman Problem (TSP), or whether a hard clause is satisfied in Boolean Satisfiability (SAT) (Cai and Lei 2020). When local search converges to a local optimum, GLS will select a subset of features presented in the incumbent solution to increase the associated penalty value, so as to force local search to explore novel solutions. In other words, GLS breaks out of the local optimum by modifying the problem’s objective landscape.
@@ 2.3 Generalized Distributed Breakout Algorithm
GDBA (Okamoto et al. 2016) provides a comprehensive set of rules for breaking out of local optima in DCOPs, which adapts the concepts of constraint violation and cost increase in DBA (Hirayama and Yokoo 2005) for solving Distributed Constraint Satisfaction Problems (DisCSPs) (Yokoo et al. 1998). Essentially, GDBA can be viewed as an instantiation of GLS where the features can be a full constraint function, a specific row or column within it, or a single cost cell, depending on the scope of changes to the penalty values during breakouts. Specifically, when a Quasi Local Minimum (QLM) (Hirayama and Yokoo 2005) is detected, each agent in QLM first identifies a set of violated constraints based on the cost values under the incumbent local solution, then independently increases the penalty associated with the features in these violated constraints by 1, which corresponds to selective penalization in GLS. Besides additive penalty like Eq. (2), GDBA’s variants also consider the multiplicative penalty:
@@ 3.1 Motivation
To analyze the dynamics of penalty values of GDBA, we consider its widely used variant $\langle M,NM,T\rangle$ on both sparse random DCOPs (RND) and meeting scheduling (MS) (Zivan et al. 2014; Maheswaran et al. 2004b) where GDBA demonstrates notably inferior and superior performance, respectively (cf. Section 4). Figure 1(a) plots the mean penalty values in all cost modifiers against iterations.
@@
Finally, agents in QLM independently increase the cost modifiers, which may cause agents to optimize different objectives. Consider a constraint $f_{ij}$ being flagged as violated for the first time. If agent $i$ is in QLM while agent $j$ doesn’t, then the cost modifier for $f_{ij}$ becomes 1 from agent $i$’s side but remains 0 from agent $j$’s side after breakout. Such mismatch introduces asymmetry in cost structures and potentially breaks the correspondence between pure strategy Nash Equilibrium and local optimum (Chapman et al. 2011; Grinshpoun et al. 2013).
@@ 3.2 DGLS Framework
Instead of using fixed rules in GDBA, our DGLS incorporates an adaptive rule for identifying violated constraints (cf. Algorithm 3). Intuitively, for each constraint $f_{ij}$ we first compute its normalized cost $\eta$, where $\check{f}_{ij}=\min_{d_{i},d_{j}}f_{ij}(d_{i},d_{j})$ and $\hat{f}_{ij}=\max_{d_{i},d_{j}}f_{ij}(d_{i},d_{j})$. Then we stochastically mark the constraint as violated with a probability of $\eta$. This also generalizes the utility-based penalization in classical GLS (Voudouris et al. 2010), where the features with the highest score are deterministically penalized.
Our adaptive rule has several unique advantages. First, it eliminates the need to tune the constraint violation condition in GDBA (e.g., NZ,NMNZ,\;NM and $MX$). Second, it perfectly aligns with objective (1) by measuring the “badness” of a constraint with the normalized cost. Particularly, if the constraint cost equals the minimum value, then the constraint cannot be flagged as violated; conversely, if the constraint cost attains the maximum value, then there must be a constraint violation. By selectively penalizing constraints, we avoid over-penalizing the constraints with the cost close to their minimum value, which directs local search to pay more attention to the constraints that incur high costs.
@@ 3.3 Theoretical Results
At the beginning of each round, the cost modifier of the constraint between agent $i$ and agent $j$ from $i$’s side is the same as the counterpart from $j$’s side. [^1] Proof is provided in the extended version.
@@
$\begin{aligned} &\textsc{EffCost}_{A}^{(k)}(d_{i},j,d_{j})=1+M_{ij}(d_{i},d_{j})\\ &\textsc{EffCost}_{M}^{(k)}(d_{i},j,d_{j})=1\cdot\left[1+M_{ij}(d_{i},d_{j})\right]=1+M_{ij}(d_{i},d_{j}).\end{aligned}$
@@
$(A,\gamma,tab)$ is equivalent to Maximum Gain Message (MGM) for any $\gamma$.
@@
This exactly matches the decision rule for agents in MGM (Maheswaran et al. 2004a). Besides,
$\begin{aligned} \Delta_{i}&=\sum_{j\in\mathcal{N}_{i}}\textnormal{{EffCost}}(d_{i},j,d_{j})-\textnormal{{EffCost}}(d_{i}^{*},j,d_{j})\\ &=\sum_{j\in\mathcal{N}_{i}}f_{ij}(d_{i},d_{j})+M_{ij}(d_{i},d_{j})-f_{ij}(d_{i}^{*},d_{j})-M_{ij}(d_{i}^{*},d_{j})\\ &=\sum_{j\in\mathcal{N}_{i}}f_{ij}(d_{i},d_{j})-f_{ij}(d_{i}^{*},d_{j})+c_{ij}-c_{ij}\\ &=\sum_{j\in\mathcal{N}_{i}}f_{ij}(d_{i},d_{j})-f_{ij}(d_{i}^{*},d_{j}),\end{aligned}$
@@
(a) Sparse problems
(b) Dense problems
Figure 2: Performance on random DCOPs